OpenStudy (anonymous):

A cube has a volume of 4 cm3. What is the new volume if each side length is doubled? A cube has a volume of 4 cm3. What is the new volume if each side length is doubled? @Mathematics

OpenStudy (anonymous):

12cm^3?

OpenStudy (unklerhaukus):

if the volume is 4 cm^3 then each side length is 4^{1/3} cm doubling this becomes 2 x 4^{1/3} cm cubing 2^3 + 4 = 8 +4 = 12 cm^3

OpenStudy (anonymous):

oh i think i get it thank you

OpenStudy (anonymous):

It is 32 cm^2 (2 x 4^{1/3})^3 is not 12 but 32

OpenStudy (anonymous):

Andras is right! x^3 = 4 doubling the sides means the volume is now (2x)^3 which is 8(x^3) or 8 x 4 = 32cm^3

OpenStudy (anonymous):

cant be its not one of the options

OpenStudy (anonymous):

then the question is flawed or is not written correctly here.

OpenStudy (anonymous):

length=4^1/3,thefore ,volume will be =(2*4^1/3)^3=32

OpenStudy (unklerhaukus):

i made an error in my last line ...length = 2 x 4^{1/3} cm cubing 2^3 x 4 = 8 x 4 = 32 cm^3